OpenStudy (bahrom7893):

Let N be a positive integer. Show that if a_n=b_n for n >= N, then Sum(a_n) and Sum(b_n) either both converge, or both diverge.

OpenStudy (bahrom7893):

@satellite73 can u help, plz?

OpenStudy (bahrom7893):

@imranmeah91 ?

OpenStudy (anonymous):

what class is it?

OpenStudy (bahrom7893):

Cal 2

OpenStudy (bahrom7893):

@Zarkon can u take a look at this too?

OpenStudy (zarkon):

the sum up to N-1 is a finite sum so it doesn't contribute to the convergence or divergence of the series.

OpenStudy (zarkon):

is assume \(\sum a_n\) converges then \[\sum_{k=1}^{\infty}b_n=\sum_{k=1}^{N-1}b_n+\sum_{k=N}^{\infty}b_n\] \[=\sum_{k=1}^{N-1}b_n+\sum_{k=N}^{\infty}a_n\] since \(\sum_{k=1}^{N-1}b_n\) is finite and \(\sum a_n\) converge...so does \(\sum b_n\)

OpenStudy (bahrom7893):

tnx a lot man u're a lifesaver!

OpenStudy (zarkon):

np